OpenStudy (anonymous):

Use algebra to evaluate the limit. lim(h-->0) 3(2+h)^2 -12/h (3 times 2+h squared minus 12 over h)

OpenStudy (anonymous):

Hint: expand the numerator, and cancel out \(h\) from the denominator.

OpenStudy (anonymous):

i know you have to expand i forgot how to expand 3(2+h)^2

OpenStudy (anonymous):

Use the FOIL method: \[3(\color{blue}{2}+\color{orange}{h})^2=3 \left[ \color{blue}{2} \cdot \color{blue}{2} + \color{blue}{2} \cdot \color{orange}{h} + \color{orange}{h} \cdot \color{blue}{2} +\color{orange}{h} \cdot \color{orange}{h} \right] = 3[4 + 4h + h^2]= \cdots\]look familiar?